OpenStudy (anonymous):

sketch a graph of f(x)=-4=3

OpenStudy (anonymous):

f(x)=-4+3

OpenStudy (anonymous):

f(x)=-4+3 f(x) = -1 http://www.wolframalpha.com/input/?i=plot+f%28x%29+%3D+-1

OpenStudy (anonymous):

ok i have another question

OpenStudy (anonymous):

Ok. What is it?

OpenStudy (anonymous):

in a certain year the amount A of garbage in pounds produced after t days by an average household is given by A=5.5 t Graph the equation for t>0